# Retail Central Bank Digital Currency:

### v1

@@ 4.2 Lobbying opportunities for private-sector incumbents
If CBDC is about providing a public payment option in the form of digital cash, then the development of CBDC could be viewed less as an extension of existing digital payments systems, and more as an adaptation of physical means of payment to the digital economy. In this case, it would stand to reason that businesses involved in these supply chains would be natural partners and instigators of any plan to create digital cash However, it is not clear that those businesses have had much of a voice in public conversations about CBDC to date.
@@ 4.3 Governments, institutional incentives, and the Panopticon
custodialnon-custodialptcard paymentsinterbank transfers ptpossible opportunityfor CBDC ptptcashbearer vouchers digitalphysical
@@ 5 Accounts all the time
Throughout the debate about the design of CBDC, the question of whether money would be “account-based” or “token-based” has featured prominently. Much of this debate centred around the assumption that token-based money, which can be non-custodial, is either incompatible with a digital payment system or incompatible with the existence or use of bank accounts. This serves as a false dichotomy between digital money and non-custodial money. The use of accounts is not a neutral design choice: Accounts intrinsically introduce at least some degree of accountability of the account-holder to the party maintaining the account. The real issue is about custody: It is possible for money to be both digital and non-custodial (see Figure 1). There are several lines of argumentation that might explain this misunderstanding:
@@ 6.1 Competition and Innovation
One central question is whether a particular direction for the design of digital currency could be expected to nurture greater competition and innovation than any other. Were the models for digital currency based upon custodial accounts to be adopted, then digital currency would be more akin to bank deposits than to cash, and it would be likely that regulated financial institutions, mostly consumer-facing banks, would provide services to users. This approach can be expected to protect the market structure for electronic payments of today, and we do not envisage any major differences. In contrast, were an approach to digital currency based upon tokens and non-custodial wallets to be adopted instead, then digital currency would be more akin to cash. For the purpose of retail payments, not only would banks, incumbent financial institutions, and money services businesses continue to provide essential services, but other kinds of businesses and organisations would have a place in the value chain as well, including those that provide ancillary services for the purposes of facilitating payments, providing hardware, maintaining software, offering better user experiences, and even offering custody for digital assets. The diversity in the range and types of services, as well as the competition and innovation afforded by lower barriers to entry, would enhance the efficiency and value delivered to users and consumers. Through the lens of competition, it is clear that designs oriented around adapting cash for the digital age, are preferable to those that seek to build a new back-end for custodial business models.
@@ 6.3 Property
In many cultures, the right to private property is broadly considered to be a foundational human right. Money is no different. English law, for example, includes the idea of a chose in action, which is something whose value in a transaction indexes the legal right of a particular party to transact it, and a chose in possession, which is something whose value in a transaction is intrinsic to holding it.
Cash might be a chose in action in the sense that it represents an obligation of the Bank of England, but it is also a chose in possession of its bearer. Bank deposits are different; from the perspective of the beneficial owner, they are claims on the assets of the bank, a chose in action only. Bank deposits have also replaced cash, in a growing number of countries, as the primary means by which individual persons make retail purchases, raising the question: If users cannot own their own money, then what can they own?
@@ 7 The future of money
Part and parcel with this debate is the establishment of groundwork for the desiderata that will govern the local and global retail payment infrastructure in the first instance, and which will extend to other areas of finance and value. The CBDC proposals seem to assume that all digital money must inexorably be held in accounts, or that there is no technically viable alternative to having third parties with institutional capabilities manage the money owned by natural persons. But argument on the basis of this assumption is casuistry, and it would seem to imply that direct ownership of money, and by extension anything of value that can be used as countertrade in an economic transaction, is a thing of the past. We reject such arguments. For us, the best model for digital currency would feature:
@@
We suggest that the core desiderata and design characteristics of public infrastructure should be agreed upon by consensus and should position the individual user at the centre, taking their interests into account as non-negotiable requirements, rather than treating them as easily curtailed pawns that the G7 so casually sets aside in furtherance of a narrow set of commercial interests.

### v2

@@ 4.2 Lobbying opportunities for private-sector incumbents
If CBDC is about providing a public payment option in the form of digital cash, then the development of CBDC could be viewed less as an extension of existing digital payments systems, and more as an adaptation of physical means of payment to the digital economy. In this case, it would stand to reason that businesses involved in the supply chains that support cash would be natural partners and instigators of any plan to create digital cash. However, it is not clear that those businesses have had much of a voice in public conversations about CBDC to date.
@@ 4.3 Governments, institutional incentives, and the Panopticon
custodialnon-custodialcard paymentsinterbank transfers possible opportunityfor CBDC cashbearer vouchers digitalphysical
@@ 5 Accounts all the time
Throughout the debate about the design of CBDC, the question of whether money would be “account-based” or “token-based” has featured prominently. Much of this debate has centred around the assumption that token-based money, which can be non-custodial, is either incompatible with a digital payment system or incompatible with the existence or use of bank accounts. This serves as a false dichotomy between digital money and non-custodial money. The use of accounts is not a neutral design choice: Accounts intrinsically introduce at least some degree of accountability of the account-holder to the party maintaining the account. The real issue is about custody: It is possible for money to be both digital and non-custodial (see Figure 1). There are several lines of argumentation that might explain this misunderstanding:
@@
  • Maintaining the power to exercise unilateral control. A final possibility is couched in the twin mechanisms of surveillance and censorship. By collecting volumes of personal data and related information about transactions, coupled with building broad profiles of individual persons who transact, service providers can more readily implement censorship of specific individuals and circumstances to meet particular objectives. Although the means to determine who and when to censor is not the same as the power to be able to censor a customer, consumer, or employee, the fact that customers are locked into accounts maintained by specific third-party actors means that the ability for customers to access and control their own assets, or to conduct transfers or transactions, can be limited and restricted, at any time and for any reason, by those third-party actors. Since custodians (and therefore governments) already enjoy such power over card and interbank payments, relinquishing that power might require an act of great courage.
@@ 6.1 Competition and Innovation
One central question is whether a particular direction for the design of digital currency could be expected to nurture greater competition and innovation than any other. Were the models for digital currency based upon custodial accounts to be adopted, then digital currency would be more akin to bank deposits than to cash, and it would be likely that regulated financial institutions, mostly consumer-facing banks, would be the only or primary entities to provide services to users, even when there is no technical reason for which this must be the case. This approach can be expected to protect the market structure for electronic payments of today, and we do not envisage any major differences. In contrast, were an approach to digital currency based upon tokens and non-custodial wallets to be adopted instead, then digital currency would be more akin to cash. For the purpose of retail payments, not only would banks, incumbent financial institutions, and money services businesses continue to provide essential services, but other kinds of businesses and organisations would have a place in the value chain as well, including those that provide ancillary services for the purposes of facilitating payments, providing hardware, maintaining software, offering better user experiences, and even offering custody for digital assets, perhaps coupled with security and management services. The diversity in the range and types of services, as well as the competition and innovation afforded by lower barriers to entry, would enhance the efficiency and value delivered to users and consumers. Through the lens of competition, it is clear that designs oriented around adapting cash for the digital age, are preferable to those that seek to build a new back-end for custodial business models. An approach based on tokens and non-custodial wallets would offer a greater diversity in the kinds and types of organisations that can service the interface between citizens and the digital economy. It would also support a way for people to choose between service providers on an as-needed basis, without being locked into account-based relationships that can be used as an instrument of surveillance and control, thus strengthening the potential for competition and thus innovation in servicing retail payments in the economy.
@@ 6.3 Property
In many cultures, the right to private property is broadly considered to be a foundational human right. Money is no different. English law, for example, includes the idea of a chose in action (a debt), which is something whose value in a transaction indexes the legal right of a particular party to transact it, and a chose in possession, which is something whose value in a transaction is intrinsic to holding it.
Cash might be a chose in action in the sense that it represents an obligation of the Bank of England, but it is also a chose in possession of its bearer as an instrument that can be held, owned, and possessed. Bank deposits are different; from the perspective of the beneficial owner, they are claims on the assets of the bank, a chose in action only. Bank deposits have also replaced cash, in a growing number of countries, as the primary means by which individual persons make retail purchases, raising the question: If users cannot own their own money, then what can they own?
@@
6.4 Addressing shortcomings of existing retail payments infrastructure
Card payment platforms dominate the existing digital payments infrastructure for retail payments. As public infrastructure, these payment platforms have important shortcomings. For example, we know that the existence of merchant fees means that merchants who charge the same price for cash payments as for card payments implicitly allow customers who pay with cash to subsidise those who pay with cards. Although the merchant fees might apply only to the card transactions, those fees amount to an economic tax on the transaction and ultimately increase the equilibrium price for consumers. Additionally, the benefits from rewards programmes offered by card schemes accrue differently to different classes of consumers. As researchers at the Federal Reserve Bank of Boston observed, “[b]ecause credit card spending and rewards are positively correlated with household income, the payment instrument transfer also induces a regressive transfer from low-income to high-income households in general” [130].
In addition to intrinsic flaws with card payments, there are extrinsic problems as well. The largest card platforms, including Visa and Mastercard, are based in the US, much to the chagrin of central banks and payment system regulators in other countries. Given the burgeoning systemic indispensability of card payments to many economies around the world, it would be understandable for businesses and governments of countries outside the US to seek strategic autonomy, if for no other reason than to protect themselves from the possibility that the US government might act adversely, for example, to introduce tariffs, or to force card platforms to withdraw from certain markets entirely.
For all of these reasons, it might be reasonable to ask whether card payments, a twentieth century invention, remain fit for purpose in the twenty-first century and beyond. Central bank digital currency might offer a way to rethink the role of card payments in the economy in general, as well as to provide a public payment option for consumers in the digital economy in particular.
@@ 7 The future of money
Part and parcel with this debate is the establishment of groundwork for the desiderata that will govern the local and global retail payment infrastructure in the first instance, and which will extend to other areas of finance and value as they digitise, broaden, and become more integrated. The CBDC proposals seem to assume that all digital money must inexorably be held in accounts, or that there is no technically viable alternative to having third parties with institutional capabilities manage the money owned by natural persons. An argument on the basis of this assumption is casuistry, and it would seem to imply that direct ownership of money, and by extension anything of value that can be used as countertrade in an economic transaction, is a thing of the past. We reject such arguments. For us, the best model for digital currency would feature, in addition to our four desiderata, the following fundamental characteristics:
@@
We suggest that the core desiderata and design characteristics of public infrastructure should be agreed upon by consensus and should position the individual user at the centre, taking their interests into account as non-negotiable requirements and foundational constraints, rather than treating them as easily curtailed pawns that the G7 and others have so casually sets aside in furtherance of a narrow set of interests.